OpenStudy (lynfran):

What is the probability of getting the number 6 at least once in a regular die if, I can roll it 6 times?

jhonyy9 (jhonyy9):

in case of one roll there is 1/6 chance to get the number of 6 yes ?

OpenStudy (lynfran):

agree

jhonyy9 (jhonyy9):

ok. so in case that can rolling 6 times so this mean that will get the chance 6*(1/6) = 6/6 = 1

OpenStudy (mathmate):

@LynFran Hint: I suggest you calculate the probability of NOT getting a 6 for six throws in a row. Subtract this probability from one to get the probability of getting a six at lease once.

OpenStudy (lynfran):

i end up with 0.665 p(y=0) ; 6C0*(1/6)^0 * (5/6)^6=0.33489 1-0.33489=0.665

satellite73 (satellite73):

doesn't seem likely

satellite73 (satellite73):

at least once mean not never

satellite73 (satellite73):

never would be \[\left(\frac{5}{6}\right)^6\]

satellite73 (satellite73):

oh whoops, you are right, sorry

satellite73 (satellite73):

\[1-(\frac{5}{6})^6=.665\]

OpenStudy (lynfran):

thanks guys